Assignment 5. Rotational dynamics Adaptive Follow-Up Problem 7.67 Constants Part A A 15 kg ball and a 3.5 kg ball are connected by a 3.0 m long rigid, massless rod. The rod and balls are rotating clockwise about its center of gravity at 22 rpm What torque will bring the balls to a halt in 7.0 s? Express your answer using two significant figures N- m Submit x Incorrect; Try Again; 4 attempts remaining Provide Feedback
Answer #1

Angular speed wi = 22 rpm = 22*2pi/60 = 2.303 rad/s

final angular speed wf = 0


angular acceleration = (0-2.303)/7 = -0.33 rad/s^2


m1 = 1.5 kg, m2 = 3.5 kg, d = 3 m

center of mass from m1, r1 = m2*d/(m1+m2) = (3.5*3)/(1.5+3.5) = 2.1 m


moment of inertia I = m1*r1^2 + m2*r2^2

r2 = d - r1 = 3 - 2.1 = 0.9 m

I = (1.5*2.1^2) + (3.5*0.9^2)

I = 9.04 kg m^2


net torque = I*alpha

torque = 9.04*0.33 = 2.98 Nm